OpenStudy (anonymous):

The density of an object is calculated using the formula D=V/M, where M is the object's mass and V is its volume. Gold has a density of 19.3 g/cm^3. What is the volume of an amount of gold that has a mass of 96.5 g?

OpenStudy (joannablackwelder):

Substitue the values into the formula.

OpenStudy (joannablackwelder):

Since the units of mass cancel, you don't have to do any conversions.

OpenStudy (joannablackwelder):

Hold on a sec. Your equation is wrong....

OpenStudy (joannablackwelder):

Density=mass/volume

OpenStudy (anonymous):

Oh yeah! I meant that

OpenStudy (joannablackwelder):

Ok,just checking! Haha

OpenStudy (joannablackwelder):

19.3=96.5/V

OpenStudy (joannablackwelder):

Can you solve for V?

OpenStudy (anonymous):

Do I multiply 96.5 to both sides?

OpenStudy (joannablackwelder):

Multiplying by 96.5 makes 96.5*19.3=96.5*96.5/V

OpenStudy (joannablackwelder):

Does that help us solve?

OpenStudy (anonymous):

So the answer will be 1,862

OpenStudy (joannablackwelder):

That 's not what I get.

OpenStudy (joannablackwelder):

To solve the equation you need to first multiply by V to get the variable in the numerator.

OpenStudy (joannablackwelder):

You can't solve for a variable as long as it is in the denominator.

OpenStudy (joannablackwelder):

That gets us to 19.3V=96.5

OpenStudy (joannablackwelder):

Make sense?

OpenStudy (joannablackwelder):

@annisasdfghjkl ?

OpenStudy (anonymous):

So multiply V where?

OpenStudy (joannablackwelder):

I should have said multiply V to both sides.

OpenStudy (wolf1728):

D = M / V 19.3 g/cm^3 = 96.5 g / V V = 96.5 g / 19.3 g / cm^3 V= ?

OpenStudy (anonymous):

Ohh okay

OpenStudy (anonymous):

After multiplying, I divide 19.3 to both sides?

OpenStudy (joannablackwelder):

Right! :)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (wolf1728):

u r welcome

OpenStudy (joannablackwelder):

No worries :)

OpenStudy (joannablackwelder):

And your unit is in cm ^3

OpenStudy (joannablackwelder):

Which comes from the density

OpenStudy (anonymous):

Okay

OpenStudy (wolf1728):

or to put it neater cm³